# Contour Flow Constraint: Preserving Global Shape Similarity for Deep Learning based Image Segmentation

### v1

@@ Abstract
For effective image segmentation, it is crucial to employ constraints informed by prior knowledge about the characteristics of the areas to be segmented to yield favorable segmentation outcomes. However, the existing methods have primarily focused on priors of specific properties or shapes, lacking consideration of the general global shape similarity from a Contour Flow (CF) perspective. Furthermore, naturally integrating this contour flow prior image segmentation model into the activation functions of deep convolutional networks through mathematical methods is currently unexplored. In this paper, we establish a concept of global shape similarity based on the premise that two shapes exhibit comparable contours. Furthermore, we mathematically derive a contour flow constraint that ensures the preservation of global shape similarity. We propose two implementations to integrate the constraint with deep neural networks. Firstly, the constraint is converted to a shape loss, which can be seamlessly incorporated into the training phase for any learning-based segmentation framework. Secondly, we add the constraint into a variational segmentation model and derive its iterative schemes for solution. The scheme is then unrolled to get the architecture of the proposed CFSSnet. Validation experiments on diverse datasets are conducted on classic benchmark deep network segmentation models. The results indicate a great improvement in segmentation accuracy and shape similarity for the proposed shape loss, showcasing the general adaptability of the proposed loss term regardless of specific network architectures. CFSSnet shows robustness in segmenting noise-contaminated images, and inherent capability to preserve global shape similarity.
@@ 1 Introduction
Traditionally, this problem can be tackled by learning-based methods and model-based methods. Learning-based methods are data-driven. They utilize parameterized deep neural networks to learn the image features and classify foreground region pixels according to the ground truth. Fully Convolutional Network (FCN)(Long et al., 2015), U-net(Ronneberger et al., 2015), DeepLab(Chen et al., 2017), and Vision-Transformer(Dosovitskiy et al., 2020) are popular learning-based segmentation frameworks with powerful learning and feature extraction ability. More recently, Segment Anything Model (SAM)(Kirillov et al., 2023) with image encoder, prompt encoder, and mask decoder architecture has achieved awesome segmentation performances. In contrast, model-based methods, such as the Potts model(Potts, 1952), Threshold Dynamics(Merriman et al., 1994), Active Contours(Kass et al., 1988), Chan-Vese model(Chan and Vese, 2001), employ variational models to minimize some energy associated with segmentation, so that the result segmentation encompasses a great fidelity. Importantly, specially designed regularization and prior-knowledge constraints could be added to the variational model to introduce desirable properties into the segmentation. Specifically, these priors include details of the area and shape of the segmented regions. (Veksler, 2008) proposed a prior to constrain the segmentation region to be a star shape. (Royer et al., 2016; Yan et al., 2020) both proposed a convexity shape constraint to segment convex regions based on Potts model(Potts, 1952) and Chan-Vese model(Chan and Vese, 2001) respectively. (Vicente et al., 2008) raised a prior to ensure topology connectivity of the segmentation region. (Rousson and Paragios, 2002) utilized the level-set representation as a shape prior so that the segmented region exhibits global consistency with a prior shape. In recent years, there has been a trend towards hybrid approaches that combine the advantages of both learning-based and model-based methods, so that the segmentation framework could leverage the power of data-driven learning while incorporating prior knowledge or constraints to improve segmentation accuracy. In(Liu et al., 2022), the star shape prior was successfully integrated into deep convolution neural network (DCNN). In the same paper, the authors also proposed a volume-preserving prior to constrain the area of the segmented region. The authors in (Shit et al., 2021) considered the topology of tabular, network-like regions and introduced a topology-preserving loss based on the structure prior.
Therefore, we aim for the shape features extracted by the encoder to be mathematically interpretable, and for the regularization of these shape features to not only be effective roughly across the entire dataset but also to have an effect on each individual sample. To this end, in this paper, we mathematically derive the contour constraint condition that ensures the global shape similarity between the segmentation and the ground truth. Meanwhile, we propose two types of implementation to integrate the proposed shape constraint with deep neural networks, enabling them to maintain the shape characteristics and preserve global shape similarity. Specifically, we define a global shape similarity by two shapes sharing similar contours and equivalently convert such similarity into a constraint by the orthogonality of segmentation’s gradient field and the ground truth contours’ tangent field. This constraint is called the contour flow (CF) constraint to achieve global shape similarity. Please refer to Figure 1 for an intuition of our idea. Then, we implement the CF constraint by two methods: (1) combine it into networks’ training process to propose the regularized shape loss, which can be seamlessly integrated into any segmentation frameworks; (2) add it into a variational segmentation model to ensure global shape similarity property of its solution segmentation, then unroll the iterative schemes for the variational model to directly obtain deep network architectures, thereby proposed the unrolled CFSSnet, which can inherently capture shape characteristics within input image, and preserve global shape similarity.
@@
2 The related work
@@ 2.1 General Variational Segmentation Framework
An image can be represented as a mapping $I:\Omega\mapsto\mathbbm{R}^{d}$, where $d$ is the image channels. Image segmentation aims to find a segmentation function $u:\Omega\mapsto\{0,1\}$ that accurately divides $\Omega$ into foreground $\{i\in\Omega|u(i)=1\}$ and background $\{i\in\Omega|u(i)=0\}$ regions. We will utilize the variational segmentation framework(Liu et al., 2022), which bridges deep neural networks and variational optimization. As proposed in(Liu et al., 2022), we call this variational model as the STD framework:
@@
where $\mathbbm{U}\coloneqq\{u|u\in C^{1}(\Omega),u:\Omega\mapsto[0,1]\}$ is the feasible domain for segmentation function. $o:\Omega\mapsto\mathbbm{R}$ is a segmentation feature function, where higher $o(i)$ indicates pixel $i$ is more likely belonging to foreground. The inner product is defined as:
@@
2.2 Unrolling
3 The Proposed Global Shape Similarity Preserving Constraint
3.1 The Signed Distance Function for an Object
Let $\Omega_{g}\coloneqq\{i\in\Omega|g(i)=1\}$ stand for an object. it can be equivalently defined by a signed distance function $\phi_{g}$ derived from its boundary $\partial\Omega_{g}\coloneqq\{i\in\Omega_{g}|\forall\mathcal{N}(i),\exists j\notin\Omega_{g}\}$, where $\mathcal{N}(i)$ stands for the adjacent pixels in $i$’s neighborhood.
@@ 3.3 Contour Flow Constraint
as the CF (Contour Flow) constraint for shape-preserving image segmentation. We will explore how to implement it into the segmentation process in the next sections. Specifically, we will propose two ways: (1) convert it to a loss function to add into any segmentation frameworks, (2) unroll a variational problem for segmentation, whose solution satisfies contour flow constraint condition, to generate a deep neural network preserving global shape similarity.
@@ 4 Implementation I: Shape Loss for any Deep-learning based Segmentation Framework
$\mathcal{L}_{Dice}$ $=1-\frac{2\sum_{i\in\Omega}u(i)g(i)}{\sum_{i\in\Omega}u^{2}(i)+\sum_{i\in\Omega}g^{2}(i)}.$ (7)
@@ 5 Implementation II: Deep Neural Network by Unrolling Variational Model
In this section, we will add the CF constraint into a general variational segmentation model, i.e. the STD framework, so that the output of DCNN will guarantee global shape similarity. Then, we will unroll the iteration schemes for the solution to generate a shape-preserving deep neural network.
@@ 5.1 Variational Segmentation Model with Global Shape Similarity
According to property of STD framework(Liu et al., 2022), it has closed form solution, $u^{t+1}=\mathcal{S}((o-{\rm div}(q^{t}\boldsymbol{F}_{g}))/\varepsilon)$, where $\mathcal{S}(\cdot)$ is sigmoid function. The $q$ sub-problem (Eq.12b) is:
@@
We solve the $q$ sub-problem by the gradient ascend method: $q^{t+1}=q^{t}-\tau(\nabla u^{t+1}\cdot\boldsymbol{F}_{g})$, where $\tau>0$ is a update rate. As analyzed above, we actually require orthogonality between $\nabla u,\boldsymbol{F}_{g}$.
Consequently, we have the following alternating iterative schemes solving variational segmentation model with global shape similarity (Eq.11), and corresponding Algorithm 1:
@@
       $q^{t+1}=q^{t}-\tau\displaystyle{\langle\nabla u^{t+1}},\boldsymbol{F}_{g}\rangle$;
@@ 5.2 CFSSnet – Unrolled DCNN Preserving Global Shape Similarity
We are now going to unroll the alternating iterative schemes for our proposed variational segmentation model with global shape similarity to obtain neural network architectures and propose the CFSSnet (Contour Flow Shape Similarity Networks) for the segmentation task. The whole architecture consists of three parts: backbone as feature extractor, two sub-branches to predict segmentation feature and contour flow and shape constraint modules to preserve shape similarity. Refer to Figure 4 for the whole architecture and components
@@ 5.2.3 Shape Constraint Module
The shape constraint module is obtained by unrolling the alternating iterative schemes. The $\nabla,{\rm div}$ operators are replaced by convolution layers with fixed kernels representing discrete gradient and divergence operators. The sigmoid function $\mathcal{S}$ corresponds to the sigmoid activation layer. According to Algorithm 1, we use $T$ shape constraint modules in total. Each module outputs an intermediate segmentation function $u^{t}$, which will be involved in computing $u^{t+1}$ of the next module.
@@ 5.2.4 Loss Function
The proposed CFSSnet has two parts of loss functions. For the first part, our network predicts a contour flow $\boldsymbol{F}$ to constrain the segmentation shape, while the true contour flow is $\boldsymbol{F}_{g}$. We use a $L^{2}$ distance to measure the loss.
@@
${\color[rgb]{0,0,0}\mathcal{L}_{CFSS}=\alpha\mathcal{L}_{flow}+\beta\mathcal{L}_{seg},\ \alpha,\beta>0.}$ (16)
@@ 6.1 Datasets and Preprocessing
  • Gastrointestinal Endoscopy Dataset (Kvasir)(Jha et al., 2020): This dataset is from the 26th international conference of MultiMedia Modeling (MMM 2020), containing 1001 endoscopy images, and is used for segmenting polyps from gastrointestinal images.
@@ 6.2 Evaluation Metrics
6.2.1 Intersection Over Union (IOU)
Let $|\Omega_{u}|$ represent the number of pixels contained in $\Omega_{u}$. The Intersection Over Union (IOU) between $\Omega_{u}$ and $\Omega_{g}$ is defined as:
${\rm\textbf{{IOU}}}(\Omega_{u},\Omega_{g})=\frac{2*|\Omega_{u}\cap\Omega_{g}|}{|\Omega_{u}|+|\Omega_{g}|}\times 100\%.$ (17)
IOU indicates the extent of overlap between the segmentation shape $\Omega_{u}$ and the ground truth. A larger value suggests that the segmentation shape $\Omega_{u}$ is closer to the ground truth.
@@ 6.2.2 Boundary Distance (BD)
where $D(i,\partial\Omega_{g})=\mathop{\min}_{j\in\partial\Omega_{g}}\|i-j\|_{2}$, i.e. the distance from $i$ to the ground truth segmentation boundary $\partial\Omega_{g}$.
This metric reflects the distance between the segmentation boundary and the ground truth boundary. A smaller distance indicates that the boundaries are closer, implying more accurate results. Additionally, if there are isolated points and salt-and-pepper noise in the segmentation shape, the boundary distance will significantly increase, making this metric a characterization of the segmentation result’s piece-wise constant smoothness.
@@ 6.2.3 Boundary Distance Standard Deviation (BDSD)
A smaller value of this metric indicates that the distances from the pixels on the segmentation boundary $\partial\Omega_{u}$ to the ground truth boundary $\partial\Omega_{g}$ exhibit less variation, suggesting a higher level of shape similarity between the two.

### v2

@@ Abstract
For effective image segmentation, it is crucial to employ constraints informed by prior knowledge about the characteristics of the areas to be segmented to yield favorable segmentation outcomes. However, the existing methods have primarily focused on priors of specific properties or shapes, lacking consideration of the general global shape similarity from a Contour Flow perspective. Furthermore, naturally integrating this contour flow prior image segmentation model into the activation functions of deep convolutional networks through mathematical methods is currently unexplored. In this paper, we establish a concept of global shape similarity based on the premise that two shapes exhibit comparable contours. Furthermore, we mathematically derive a contour flow constraint that ensures the preservation of global shape similarity. We propose two implementations to integrate the constraint with deep neural networks. Firstly, the constraint is converted to a shape loss, which can be seamlessly incorporated into the training phase for any learning-based segmentation framework. Secondly, we add the constraint into a variational segmentation model and derive its iterative schemes for solution. The scheme is then unrolled to get the architecture of the proposed CFSSnet. Validation experiments on diverse datasets are conducted on classic benchmark deep network segmentation models. The results indicate a great improvement in segmentation accuracy and shape similarity for the proposed shape loss, showcasing the general adaptability of the proposed loss term regardless of specific network architectures. CFSSnet shows robustness in segmenting noise-contaminated images, and inherent capability to preserve global shape similarity.
Keywords: Image segmentation; DCNN; global shape similarity; contour flow constraint; shape loss; unrolled network
@@ 1 Introduction
In this paper, we mathematically derive the contour constraint condition that ensures the global shape similarity between the segmentation and the ground truth. Meanwhile, we propose two types of implementation to integrate the proposed shape constraint with deep neural networks, enabling them to maintain the shape characteristics and preserve global shape similarity. Specifically, we define a global shape similarity by two shapes sharing similar contours and equivalently convert such similarity into a constraint by the orthogonality of segmentation’s gradient field and the ground truth contours’ tangent field. This constraint is called the contour flow (CF) constraint to achieve global shape similarity. Please refer to Figure 1 for an intuition of our idea. Then, we implement the CF constraint by two methods: (1) combine it into networks’ training process to propose the regularized shape loss, which can be seamlessly integrated into any segmentation frameworks; (2) add it into a variational segmentation model to ensure global shape similarity property of its solution segmentation, then unroll the iterative schemes for the variational model to directly obtain deep network architectures, thereby proposed the unrolled CFSSnet, which can inherently capture shape characteristics within input image, and preserve global shape similarity.
@@
2 Related work
2.1 Learning Based and Model Based Segmentation
Traditionally, image segmentation can be tackled by learning-based methods and model-based methods. Learning-based methods are data-driven. They utilize parameterized deep neural networks to learn the image features and classify foreground region pixels according to the ground truth. Fully Convolutional Network (FCN) (Long et al., 2015), U-net (Ronneberger et al., 2015), DeepLab (Chen et al., 2017), and Vision-Transformer (Dosovitskiy et al., 2020) are popular learning-based segmentation frameworks with powerful learning and feature extraction ability. More recently, Segment Anything Model (SAM) (Kirillov et al., 2023) with image encoder, prompt encoder, and mask decoder architecture has achieved awesome segmentation performances. In contrast, model-based methods, such as the Potts model (Potts, 1952), Threshold Dynamics (Merriman et al., 1994), Active Contours (Kass et al., 1988), Chan-Vese model (Chan and Vese, 2001), employ variational models to minimize some energy associated with segmentation, so that the result segmentation encompasses a great fidelity. Importantly, specially designed regularization and prior-knowledge constraints could be added to the variational model to introduce desirable properties into the segmentation. Specifically, these priors include details of the area and shape of the segmented regions. (Veksler, 2008) proposed a prior to constrain the segmentation region to be a star shape. (Royer et al., 2016; Yan et al., 2020) both proposed a convexity shape constraint to segment convex regions based on Potts model (Potts, 1952) and Chan-Vese model (Chan and Vese, 2001) respectively. (Vicente et al., 2008) raised a prior to ensure topology connectivity of the segmentation region.  (Rousson and Paragios, 2002) utilized the level-set representation as a shape prior so that the segmented region exhibits global consistency with a prior shape. In recent years, there has been a trend towards hybrid approaches that combine the advantages of both learning-based and model-based methods, so that the segmentation framework could leverage the power of data-driven learning while incorporating prior knowledge or constraints to improve segmentation accuracy. In (Liu et al., 2022), the star shape prior was successfully integrated into deep convolution neural network (DCNN). In the same paper, the authors also proposed a volume-preserving prior to constrain the area of the segmented region. The authors in (Shit et al., 2021) considered the topology of tabular, network-like regions and introduced a topology-preserving loss based on the structure prior.
2.2 Shape Prior in Image Segmentation
Therefore, we aim for the shape features extracted by the encoder to be mathematically interpretable, and for the regularization of these shape features to not only be effective roughly across the entire dataset but also to have an effect on each individual sample. This can be achieved by our proposed contour flow constraint with loss function and variational unrolling implementation.
2.3 Unrolling
3 The Proposed Global Shape Similarity Preserving Constraint
@@ 3.1 General Variational Segmentation Framework
An image can be represented as a mapping $I:\Omega\mapsto\mathbbm{R}^{d}$, where $d$ is the image channels. Image segmentation aims to find a segmentation function $u:\Omega\mapsto\{0,1\}$ that accurately divides $\Omega$ into foreground $\{i\in\Omega|u(i)=1\}$ and background $\{i\in\Omega|u(i)=0\}$ regions. We will utilize the variational segmentation framework (Liu et al., 2022), which bridges deep neural networks and variational optimization. As proposed in (Liu et al., 2022), we call this variational model as the soft threshold framework:
@@
where $\mathbbm{U}\coloneqq\{u|u\in C^{1}(\Omega),u:\Omega\mapsto[0,1]\}$ is the feasible domain for segmentation function. $o:\Omega\mapsto\mathbbm{R}$ is a segmentation feature function, where the higher $o(i)$ indicates that the pixel $i$ is more likely to belong to the foreground. The inner product is defined as:
@@
3.2 The Signed Distance Function for Ground Truth
Let $\Omega_{g}\coloneqq\{i\in\Omega|g(i)=1\}$ stand for ground truth, it can be equivalently defined by a signed distance function $\phi_{g}$ derived from its boundary $\partial\Omega_{g}\coloneqq\{i\in\Omega_{g}|\forall\mathcal{N}(i),\exists j\notin\Omega_{g}\}$, where $\mathcal{N}(i)$ stands for the adjacent pixels in $i$’s neighborhood.
@@ 3.4 Contour Flow Constraint
as the Contour Flow (CF) constraint for shape-preserving image segmentation. We will explore how to implement it into the segmentation process in the next sections. Specifically, we will propose two ways: (1) convert it to a loss function to add into any segmentation frameworks, (2) unroll a variational problem for segmentation, whose solution satisfies contour flow constraint condition, to generate a deep neural network preserving global shape similarity.
@@ 4 Implementation I: Shape Loss for any Deep-learning based Segmentation Framework
$\mathcal{L}_{Dice}$ $=1-\frac{2\sum_{i\in\Omega}u(i)g(i)}{\sum_{i\in\Omega}u(i)+\sum_{i\in\Omega}g(i)}.$ (7)
@@ 5 Implementation II: Deep Neural Network by Unrolling Variational Model
In this section, we will add the CF constraint into a general variational segmentation model, i.e. the soft threshold classification framework, so that the output of DCNN will guarantee global shape similarity. Then, we will unroll the iteration schemes for the solution to generate a shape-preserving deep neural network.
@@ 5.1 Variational Segmentation Model with Global Shape Similarity
According to property of framework (Liu et al., 2022), it has closed form solution, $u^{t+1}=\mathcal{S}((o-{\rm div}(q^{t}\boldsymbol{F}_{g}))/\varepsilon)$, where $\mathcal{S}(\cdot)$ is sigmoid function. The $q$ sub-problem (Eq.12b) is:
@@
We solve the $q$ sub-problem by the gradient ascend method: $q^{t+1}=q^{t}-\tau(\nabla u^{t+1}\cdot\boldsymbol{F}_{g})$, where $\tau>0$ is a update rate. As analyzed above, we actually require orthogonality between $\nabla u,\boldsymbol{F}_{g}$. The $q$ variable here determines the bias term to be reduced on feature $o$ according to the extent of “how orthogonal $\nabla u,\boldsymbol{F}_{g}$ are”. When orthogonality is met, $q$ reaches optimal and stops to update. Therefore, at early steps, larger $\tau$ helps to converge to optimal value of $q$ more quickly. However, when $\nabla u\cdot\boldsymbol{F}_{g}$ gets small, large $\tau$ may cause $q$ to “overshoot” and swing around the optimal. Empirically, setting $\tau=10$ can get a trade-off between speed and accuracy.
Consequently, we have the following schemes to solve the variational segmentation model with global shape similarity (Eq.11), and the corresponding Algorithm 1:
@@
       $q^{t+1}=q^{t}-\tau\displaystyle{(\nabla u^{t+1}}\cdot\boldsymbol{F}_{g})$;
@@ 5.2 CFSSnet – Unrolled DCNN Preserving Global Shape Similarity
We are now going to unroll the alternating iterative schemes for our proposed variational segmentation model with global shape similarity to obtain neural network architectures and propose the CFSSnet (Contour Flow Shape Similarity Networks) for the segmentation task. The whole architecture consists of three parts: backbone as feature extractor, two sub-branches to predict segmentation feature and contour flow and shape constraint modules to preserve shape similarity. Refer to Figure 4 for the whole architecture and components.
@@ 5.2.3 Shape Constraint Module
The shape constraint module is obtained by unrolling the alternating iterative schemes. The $\nabla,{\rm div}$ operators are implemented by convolution layers with fixed kernels of finite differences stencils representing discrete gradient and divergence operators. We use first-order scheme $\partial_{x}u=u(i_{x}+1,i_{y})-u(i_{x},i_{y})$, and $\partial_{y}u=u(i_{x},i_{y}+1)-u(i_{x},i_{y})$. The sigmoid function $\mathcal{S}$ corresponds to the sigmoid activation layer. According to Algorithm 1, we use $T$ shape constraint modules in total. Each module outputs an intermediate segmentation function $u^{t}$, which will be involved in computing $u^{t+1}$ of the next module.
@@ 5.2.4 Loss Function
The proposed CFSSnet has two parts of loss functions. For the first part, our network predicts a contour flow $\boldsymbol{F}$ for the segmentation shape, resembling true contour flow $\boldsymbol{F}_{g}$. We use a $L^{2}$ loss to guide this learning process:
@@
Although learning the contour flow $\boldsymbol{F}_{g}$ encodes higher-order (derivative) information that may seem more challenging than predicting the segmentation map itself, its main function is to impose a neighborhood-based regularization on the soft-classification $u$. In a standard softmax segmentation network, each pixel is classified independently using only its local features $o$. Such an approach ignores how adjacent pixels should cohere. By introducing the additional CF prediction branch and minimizing the flow-consistency loss we force the predicted flow $\boldsymbol{F}$ to align with the true gradient $\boldsymbol{F}_{g}$, encouraging $u$ to respect its neighbors. The subsequent updates in Eqs.(13a) and (13b) use this learned flow to refine $u$ in a way that naturally incorporates pixel-neighborhood information. Consequently, even if $\boldsymbol{F}$ is not perfectly accurate, its presence still improves segmentation performance compared to a baseline that classifies pixels independently. We also have an ablation study to quantitatively investigate potential influence of flow estimation errors.
@@
$\mathcal{L}_{CFSS}=\alpha\mathcal{L}_{flow}+\beta\mathcal{L}_{seg},\ \alpha,\beta>0.$ (16)
@@ 6.1 Datasets and Preprocessing
  • Gastrointestinal Endoscopy Dataset (Kvasir) (Jha et al., 2019): This dataset is from the 26th international conference of MultiMedia Modeling (MMM 2020), containing 1001 endoscopy images, and is used for segmenting polyps from gastrointestinal images.
@@ 6.2 Evaluation Metrics
6.2.1 Dice Similarity Coefficient (Dice)
Let $|\Omega_{u}|$ represent the number of pixels contained in $\Omega_{u}$. The Dice score between $\Omega_{u}$ and $\Omega_{g}$ is defined as:
${\rm\textbf{{Dice}}}(\Omega_{u},\Omega_{g})=\frac{2*|\Omega_{u}\cap\Omega_{g}|}{|\Omega_{u}|+|\Omega_{g}|}\times 100\%.$ (17)
Dice indicates the extent of overlap between the segmentation shape $\Omega_{u}$ and the ground truth. A larger value suggests that the segmentation shape $\Omega_{u}$ is closer to the ground truth.
@@ 6.2.2 Boundary Distance (BD)
where $D(i,\partial\Omega_{g})=\mathop{\min}_{j\in\partial\Omega_{g}}\|i-j\|_{2}$, i.e. the distance from $i$ to the ground truth segmentation boundary $\partial\Omega_{g}$. This metric reflects the average distance from the segmentation boundary to the ground truth boundary. A smaller distance indicates that the boundaries are closer, implying more accurate results.
@@ 6.2.3 Boundary Distance Standard Deviation (BDSD)
This is the square root of variance of $D(i,\partial\Omega_{g})$. If the segmentation boundary is too distinct, even a shape with small BD could have large BDSD. A smaller value of this metric indicates that the distances from the pixels on $\partial\Omega_{u}$ to $\partial\Omega_{g}$ exhibit more consistency, suggesting a higher level of shape similarity between the two. If BDSD is equal to 00, $\Omega_{u},\Omega_{g}$ has exactly the shape similarity defined in Definition 3.